# Enabling Fine-Grained Operating Points for Black-Box LLMs

### v1

@@ 3.2 Analysis of LLM-Generated Probability Estimates
Figure 2 (a): All 3 LLMs are heavily biased towards verbalizing probability estimates that end with "0" and "5"s. On the other hand, the distributions of the 3rd characters are spread across all tokens (digits) with slights bias towards extremes. Interestingly we observe that even though all three models’ top 2 most frequent tokens in position 4 are consistent, Qwen’s highest probability token is "5" while the others’ is "0".
@@
Verbalized classification scores have low cardinality. Figure 1 depicts the PR and ROC curves of the distributions $\hat{y}^{\text{vrb}}_{\text{Claude}}$, $\hat{y}^{\text{vrb}}_{\text{Nova}}$, $\hat{y}^{\text{vrb}}_{\text{Qwen}}$ and when prompted to output class probabilities, and $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ when prompted to output binary decisions (e.g., yes or no). From the figure, it can be observed that all three LLMs generate sparsely populated PR and ROC curves when verbalized probabilities are used, while the token probabilities of Qwen (i.e., $\hat{y}_{\text{Qwen}}^{tkn}$) form continuous curves. 2Figure 2 (b): shows output cardinalities and operational granularities of these four variants. We observe that using verbalized probabilities only provide 16-23 unique values, while token probabilities provide 41671 unique values. Notably, we observe minimal change in cardinality when we prompt the LLMs to generate scores in [0, 100], generate and combine a coarse and a fine grained score, generate 20 scores within a single output and conduct post-processing to aggregate them, and generate output by using a context of 20 example scores in desired high-precision form. See Appendix A.4 for experiment results with these variants. We additionally consider two prompt based approaches from state-of-the-art confidence estimation works Tian et al. (2023); Xiong et al. (2023). These are two stage approaches (requiring $2$ LLM calls per instance) with the LLM first predicting just the class label and then provided with this label, predicting the likelihood of its correctness. One of these variants uses chain-of-thought prompting in addition to the two-stage approach.
@@
Verbalized probability generation is biased towards round but informative numbers. In order to gain additional insights into low-cardinality generations, we dive deeper into token probabilities associated with the generations $\hat{y}^{\text{vrb}}_{\text{Qwen}}$ character by character. Note that, this is different than $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ that corresponds to the token probabilities associated to binary decisions. Specifically, given a verbalized probability $\hat{y}^{\text{vrb}}[i]$, we individually analyze the token probability distribution of each character by their position. For example, let $\hat{y}^{\text{vrb}}_{1}=$"0.95". Then, the positions 1 and 3 would correspond to the tokens "0" and "9", respectively. For the black-box Claude and Nova, because the token probabilities are not available, we conduct the same analysis by counting the characters occurring at each position. 2Figure 2 (a): depicts the distribution of the token probabilities and character counts by position across 11 datasets. From the figure, it can be observed that when simply asked to verbalize class probabilities, all three LLMs chose to end their generations with "0" and "5" tokens most of the time, regardless of the task. We hypothesize that this is due to the fact that LLMs are trained and aligned by heavily relying on human-generated data, and humans tend to prefer expressing quantities in round numbers as their cognitive bias Pope & Simonsohn (2011); Zhou et al. (2023b). When this bias is reflected to LLM generations, it creates a many-to-one mapping from an internal high-precision representation (e.g., discrete token probabilities) to a round number, reducing the cardinality of generations. We refer to this generation behavior of LLMs as rounding bias. To be clear, in this context, we are using the definition of roundness as described by Ferson et al. (2015): a number ending in the digit 5 is considered rounder than any number ending in 1, 2, 3, 4, 6, 7, 8, or 9. However, a number ending in 5 is considered less round than any number ending in 0. From the figure, we also observe that the distribution in the third position is spread across all numerical tokens, while having slight bias towards extremes. Notably, even though the verbalized probabilities are heavily impacted by the rounding bias, they still imply operational curves that are visibly correlated with the ones generated using $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ as shown in Figure 1.
@@ 3.3 Proposed Approaches
where $S_{f}$ denotes the operational curve $S([y,f(\hat{y}_{i};\theta_{f})]_{i=1}^{N})$ defined for $S\in\{\mathrm{ROC},\mathrm{PR}\}$ in Equation equation 1, $g(S_{f})$ is the operational granularity of $S_{f}$ from Equation equation 2, $\hat{y}=f(\ .;\theta_{f})$ is a function that maps verbalized probabilities $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{N}$ to an output distribution $\hat{y}_{i}]_{i=1}^{N}$ that estimates class probabilities, and $\ell$ is a function that measures the classification loss. Then, the objective is to find a function $f$ (by estimating $\theta_{f}$) that minimizes the operational granularity for $a$ and $b$ of $S\in\mathbb{R}^{a\times b}$, while keeping the empirical risk bounded by the loss incurred by the original verbalized predictions $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{N}$. One way to attempt solving this optimization problem is to observe that according to Equations equation 1 and equation 2, $g^{a}(S_{f})$ and $g^{b}(S_{f})$ depend on the cardinality $\big|[\hat{y}_{i}]_{i=1}^{N}\big|$ and the entropy $H\big([\hat{y}_{i}]_{i=1}^{N}\big)$ of the output distribution. In fact, setting $\big|[\hat{y}_{i}]_{i=1}^{N}\big|$ to samples from the standard uniform distribution $U(0,1)$ would minimize this objective, however, would also violate the inequality constraint. Based on this observation, instead of directly minimizing $\big[g^{a}(S_{f}),g^{b}(S_{f})\big]$, we reformulate the objective of our first proposed variant as follows:
@@
where $\mathrm{clip}_{[0,1]}(x)$ maps $x$ to the closest value in $[0,1]$, and $z\sim{U}(0,1)$. Note that in Equation equation 4, $w$ controls the amount of noise introduced to the verbalized predictions. Then, Equation equation 4 aims to find the largest $w$ that does not impact the classification performance. Note that, to maintain the classification performance in PR and ROC spaces, it is not required to directly calculate the classification losses $\ell$. As long as the relative ranks of the predictions remain the same, the performance in these spaces will remain unchanged. Then, given a list of verbalized probabilities $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{n}$, one can set the $w$ such that ${z_{n+1}}{w}+\hat{y}_{n+1}^{\text{vrb}}<\hat{y}_{i}^{\text{vrb}}$, where $\hat{y}_{i}^{\text{vrb}}$ is the smallest probability in the list that is larger than $\hat{y}_{n+1}^{\text{vrb}}$. We refer to this approach as Ours-Unsup.
@@
where $z\sim\mathcal{N}(0,1)$ and $\mathrm{sig(x)=1/(1+e^{-x})}$. In Equation equation 5, $f$ models additive corrections made to the samples from a Gaussian distribution with the standard deviation $\sigma\propto 1/w$. Since the differential entropy of a Gaussian (i.e., $\ln(\sigma\sqrt{2\pi e}))$ is proportional to its standard deviation, decreasing the value of a positive $w$ would increase the entropy of the sample distribution that is corrected by $f$ and $z_{i}/w]_{i=1}^{N}$. Consequently, $\lambda$ controls the magnitude of regularization that pushes the entropy up, and we leave it as a hyperparameter. $(w,\theta_{f})$ is estimated using a ReLU network Hanin & Rolnick (2019) by setting $\ell$ to cross-entropy loss Mao et al. (2023). We refer to this approach as Ours-Sup-1call. It is important to note that Equation equation 5 learns a mapping from $(z_{i},\hat{y}^{\text{vrb}}_{i})$ to $y_{i}$, without having visibility of input features $\mathbf{x}_{i}$. Therefore, when $w\rightarrow\infty$, the mapping $f$ reduces to a calibrator that learns to correct $\hat{y}^{\text{vrb}}$’s mistakes that is consistently repeated across different input instances. In Appendix A.7, we provide ablation studies to show that jointly learning to introduce noise and correcting miscalibrated behavior (i.e., equation 5) performs better than conducting these steps separately.
@@ 4 Experiments
Figure 5 (a): Scatter plot of precision-recall values for all possible thresholds. The plots along the margin show the histogram of operating points for each method. The proposed methods have significantly higher output cardinality, providing better control over the operating point, while simultaneously improving the performance.
@@
Figure 5 (b): Operational granularity of the baselines and proposed approaches measured along precision (prepre), recall (recrec) and false positive rate (fprfpr) axes to cover both PR and ROC spaces, as well as their output cardinalities (carcar). Granularities are calculated using Equation equation 2. Proposed methods increase cardinality by nearly 33 orders of magnitude providing great granularity along all 33 axes. $2\sigma$- error margins are denoted in parenthesis for PRAUC.
We empirically demonstrate the improvements in performance and granularity of operating points using the proposed approaches. We use 11 publicly available binary classification datasets to evaluate the effectiveness of the proposed approach, covering a variety of applications from sentiment classification to heart disease detection. Dataset sources, sizes and processing details can be found in Appendix A.2. Similar to Section 3.2, we experiment with three LLMs of varying parameter counts - Anthropic Claude 3.0 Sonnet (Claude), Amazon Nova Pro 1.0 (Nova) and DeepSeek R1 Distill-Qwen-32B (Qwen). We implement Ours-Sup-1call as an MLP with $2$ hidden-layers of width $8$ and with ReLU activation function, and tune learning rates via grid search within $[0.01,0.05,0.1]$. We also conduct grid search to find a suitable value for $\lambda$ from Equation equation 5 within $\mathrm{LogUniform}(1e^{-4},1e^{-1})$.
@@
Results on a Simulated Diverse Dataset. Real-world scenarios often involve highly diverse data that contains sub-populations with different distributions. Due to their foundational nature, LLMs are commonly used in these scenarios. To simulate experiments on such a dataset, we combine the 11 datasets we individually experimented with into a single one and analyze the performance. For a single operating point to work well on this combined dataset, it is necessary that the predicted probabilities have consistent meaning across datasets. We observe that the supervised approaches trained on this combined dataset are particularly beneficial in such scenarios.  5Figure 5 (a): compares the precision-recall curves of different approaches for Claude. For a given approach, we use every unique value in the prediction as a threshold and calculate the corresponding precision and recall and visualize this as a scatter plot. The histograms on the margins denote the number of such points within a bin. We observe that compared to Prompt-Naive, proposed method significantly improves the operational granularity by increasing the diversity and cardinality of the outputs, while outperforming it in terms of PRAUC. In addition, we observe that even though it is significantly more efficient and scalable, Ours-Sup-2call outperforms Sample-Prob while achieving comparable or higher density in the operational curves. Unlike the individual datasets, the performance of Ours-Unsup here is not comparable to that of the supervised ones. However, it still offers an easy way to increase the cardinality while outperforming Prompt-Naive. Figure 4 (bottom) shows PRAUC comparison for the joint dataset. Unlike on the individual datasets, the proposed supervised approaches and Sample-Prob vastly outperform other methods. With just $2$ samples per instance during inference, the Ours-Sup-2call achieves $8\%$ point improvement in mean PRAUC over the baseline Prompt-Naive and $2.6\%$ points over $20$ sample Sample-Prob approach. The learning approaches have relatively higher variance due to the small sizes of dataset-subsets that are merged to create the combined dataset. 5Figure 5 (b): shows the operational granularity measurements along $g^{pre}$, $g^{rec}$ and $g^{fpr}$, as well as the output cardinalities. Prompt-Naive and the two more complex prompting techniques Prompt-Seq and Prompt-Seq-CoT have an extremely small cardinality ($16$ or less). Sampling approaches improve it by nearly two orders of magnitude. However, the proposed methods achieve a cardinality nearly equaling the size of the dataset. This increase in cardinality also translates to better operating points, seen in the low granularity values along all $3$ axes of precision, recall and fpr, while maintaining or improving performance.
@@ 5 Conclusion
We studied the impact of prompting, confidence elicitation and uncertainty estimation techniques to black-box LLMs’ output distributions for decision making problems. We showed that verbalized probabilities have limited cardinality and provided possible explanations for why and when this issue surfaces. Our proposed approaches expand the range of operating points while maintaining or improving the overall prediction quality. Although empirically successful, our approaches do not directly optimize the objective equation 3, but rely on noise to increase granularity. This results in always operating in a granularity-performance tradeoff that may fail to retain performance on tasks with highly complex relationships between $\hat{y}$ and $y$. Future work includes moving away from this tradeoff by modeling $z$ as a prior to a generative model, and fitting a classifier to the representations of this model.

### v2

@@ 3.2 Analysis of LLM-Generated Probability Estimates
Figure 2: All 3 LLMs are heavily biased towards verbalizing probability estimates that end with "0" and "5"s. On the other hand, the distributions of the 3rd characters are spread across all tokens (digits) with slights bias towards extremes. Interestingly we observe that even though all three models’ top 2 most frequent tokens in position 4 are consistent, Qwen’s highest probability token is "5" while the others’ is "0".
@@
Verbalized classification scores have low cardinality. Figure 1 depicts the PR and ROC curves of the distributions $\hat{y}^{\text{vrb}}_{\text{Claude}}$, $\hat{y}^{\text{vrb}}_{\text{Nova}}$, $\hat{y}^{\text{vrb}}_{\text{Qwen}}$ and when prompted to output class probabilities, and $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ when prompted to output binary decisions (e.g., yes or no). From the figure, it can be observed that all three LLMs generate sparsely populated PR and ROC curves when verbalized probabilities are used, while the token probabilities of Qwen (i.e., $\hat{y}_{\text{Qwen}}^{tkn}$) form continuous curves. 1 shows output cardinalities and operational granularities of these four variants. We observe that using verbalized probabilities only provide 16-23 unique values, while token probabilities provide 41671 unique values. Notably, we observe minimal change in cardinality when we prompt the LLMs to generate scores in [0, 100], generate and combine a coarse and a fine grained score, generate 20 scores within a single output and conduct post-processing to aggregate them, and generate output by using a context of 20 example scores in desired high-precision form. See Appendix A.4 for experiment results with these variants. We additionally consider two prompt based approaches from state-of-the-art confidence estimation works Tian et al. (2023); Xiong et al. (2023). These are two stage approaches (requiring $2$ LLM calls per instance) with the LLM first predicting just the class label and then provided with this label, predicting the likelihood of its correctness. One of these variants uses chain-of-thought prompting in addition to the two-stage approach.
@@
Verbalized probability generation is biased towards round but informative numbers. In order to gain additional insights into low-cardinality generations, we dive deeper into token probabilities associated with the generations $\hat{y}^{\text{vrb}}_{\text{Qwen}}$ character by character. Note that, this is different than $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ that corresponds to the token probabilities associated to binary decisions. Specifically, given a verbalized probability $\hat{y}^{\text{vrb}}[i]$, we individually analyze the token probability distribution of each character by their position. For example, let $\hat{y}^{\text{vrb}}_{1}=$"0.95". Then, the positions 1 and 3 would correspond to the tokens "0" and "9", respectively. For the black-box Claude and Nova, because the token probabilities are not available, we conduct the same analysis by counting the characters occurring at each position. Figure 2 depicts the distribution of the token probabilities and character counts by position across 11 datasets. From the figure, it can be observed that when simply asked to verbalize class probabilities, all three LLMs chose to end their generations with "0" and "5" tokens most of the time, regardless of the task. We hypothesize that this is due to the fact that LLMs are trained and aligned by heavily relying on human-generated data, and humans tend to prefer expressing quantities in round numbers as their cognitive bias Pope & Simonsohn (2011); Zhou et al. (2023b). When this bias is reflected to LLM generations, it creates a many-to-one mapping from an internal high-precision representation (e.g., discrete token probabilities) to a round number, reducing the cardinality of generations. We refer to this generation behavior of LLMs as rounding bias. To be clear, in this context, we are using the definition of roundness as described by Ferson et al. (2015): a number ending in the digit 5 is considered rounder than any number ending in 1, 2, 3, 4, 6, 7, 8, or 9. However, a number ending in 5 is considered less round than any number ending in 0. From the figure, we also observe that the distribution in the third position is spread across all numerical tokens, while having slight bias towards extremes. Notably, even though the verbalized probabilities are heavily impacted by the rounding bias, they still imply operational curves that are visibly correlated with the ones generated using $\hat{y}^{\text{tkn}}_{\text{Qwen}}$ as shown in Figure 1.
@@ 3.3 Proposed Approaches
where $S_{f}$ denotes the operational curve $S([y,f(\hat{y}_{i};\theta_{f})]_{i=1}^{N})$ defined for $S\in\{\mathrm{ROC},\mathrm{PR}\}$ in Equation 1, $g(S_{f})$ is the operational granularity of $S_{f}$ from Equation 2, $\hat{y}=f(\ .;\theta_{f})$ is a function that maps verbalized probabilities $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{N}$ to an output distribution $\hat{y}_{i}]_{i=1}^{N}$ that estimates class probabilities, and $\ell$ is a function that measures the classification loss. Then, the objective is to find a function $f$ (by estimating $\theta_{f}$) that minimizes the operational granularity for $a$ and $b$ of $S\in\mathbb{R}^{a\times b}$, while keeping the empirical risk bounded by the loss incurred by the original verbalized predictions $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{N}$. One way to attempt solving this optimization problem is to observe that according to Equation 1 and Equation 2, $g^{a}(S_{f})$ and $g^{b}(S_{f})$ depend on the cardinality $\big|[\hat{y}_{i}]_{i=1}^{N}\big|$ and the entropy $H\big([\hat{y}_{i}]_{i=1}^{N}\big)$ of the output distribution. In fact, setting $\big|[\hat{y}_{i}]_{i=1}^{N}\big|$ to samples from the standard uniform distribution $U(0,1)$ would minimize this objective, however, would also violate the inequality constraint. Based on this observation, instead of directly minimizing $\big[g^{a}(S_{f}),g^{b}(S_{f})\big]$, we reformulate the objective of our first proposed variant as follows:
@@
where $\mathrm{clip}_{[0,1]}(x)$ maps $x$ to the closest value in $[0,1]$, and $z\sim{U}(0,1)$. Note that in Equation 4, $w$ controls the amount of noise introduced to the verbalized predictions. Then, Equation 4 aims to find the largest $w$ that does not impact the classification performance. Note that, to maintain the classification performance in PR and ROC spaces, it is not required to directly calculate the classification losses $\ell$. As long as the relative ranks of the predictions remain the same, the performance in these spaces will remain unchanged. Then, given a list of verbalized probabilities $\hat{y}^{\text{vrb}}_{i}]_{i=1}^{n}$, one can set the $w$ such that ${z_{n+1}}{w}+\hat{y}_{n+1}^{\text{vrb}}<\hat{y}_{i}^{\text{vrb}}$, where $\hat{y}_{i}^{\text{vrb}}$ is the smallest probability in the list that is larger than $\hat{y}_{n+1}^{\text{vrb}}$. We refer to this approach as Ours-Unsup.
@@
where $z\sim\mathcal{N}(0,1)$ and $\mathrm{sig(x)=1/(1+e^{-x})}$. In Equation 5, $f$ models additive corrections made to the samples from a Gaussian distribution with the standard deviation $\sigma\propto 1/w$. Since the differential entropy of a Gaussian (i.e., $\ln(\sigma\sqrt{2\pi e}))$ is proportional to its standard deviation, decreasing the value of a positive $w$ would increase the entropy of the sample distribution that is corrected by $f$ and $z_{i}/w]_{i=1}^{N}$. Consequently, $\lambda$ controls the magnitude of regularization that pushes the entropy up, and we leave it as a hyperparameter. $(w,\theta_{f})$ is estimated using a ReLU network Hanin & Rolnick (2019) by setting $\ell$ to cross-entropy loss Mao et al. (2023). We refer to this approach as Ours-Sup-1call. It is important to note that Equation 5 learns a mapping from $(z_{i},\hat{y}^{\text{vrb}}_{i})$ to $y_{i}$, without having visibility of input features $\mathbf{x}_{i}$. Therefore, when $w\rightarrow\infty$, the mapping $f$ reduces to a calibrator that learns to correct $\hat{y}^{\text{vrb}}$’s mistakes that is consistently repeated across different input instances. In Appendix A.7, we provide ablation studies to show that jointly learning to introduce noise and correcting miscalibrated behavior (i.e., Equation 5) performs better than conducting these steps separately.
@@ 4 Experiments
We empirically demonstrate the improvements in performance and granularity of operating points using the proposed approaches. We use 11 publicly available binary classification datasets to evaluate the effectiveness of the proposed approach, covering a variety of applications from sentiment classification to heart disease detection. Dataset sources, sizes and processing details can be found in Appendix A.2. Similar to Section 3.2, we experiment with three LLMs of varying parameter counts - Anthropic Claude 3.0 Sonnet (Claude), Amazon Nova Pro 1.0 (Nova) and DeepSeek R1 Distill-Qwen-32B (Qwen). We implement Ours-Sup-1call as an MLP with $2$ hidden-layers of width $8$ and with ReLU activation function, and tune learning rates via grid search within $[0.01,0.05,0.1]$. We also conduct grid search to find a suitable value for $\lambda$ from Equation 5 within $\mathrm{LogUniform}(1e^{-4},1e^{-1})$.
Figure 5: Scatter plot of precision-recall values for all possible thresholds. The plots along the margin show the histogram of operating points for each method. The proposed methods have significantly higher output cardinality, providing better control over the operating point, while simultaneously improving the performance.
@@
Table 3: Operational granularity of the baselines and proposed approaches measured along precision (prepre), recall (recrec) and false positive rate (fprfpr) axes to cover both PR and ROC spaces, as well as their output cardinalities (carcar). Granularities are calculated using Equation equation 2. Proposed methods increase cardinality by nearly 33 orders of magnitude providing great granularity along all 33 axes. $2\sigma$- error margins are denoted in parenthesis for PRAUC.
@@
Results on a Simulated Diverse Dataset. Real-world scenarios often involve highly diverse data that contains sub-populations with different distributions. Due to their foundational nature, LLMs are commonly used in these scenarios. To simulate experiments on such a dataset, we combine the 11 datasets we individually experimented with into a single one and analyze the performance. For a single operating point to work well on this combined dataset, it is necessary that the predicted probabilities have consistent meaning across datasets. We observe that the supervised approaches trained on this combined dataset are particularly beneficial in such scenarios. Figure 5 compares the precision-recall curves of different approaches for Claude. For a given approach, we use every unique value in the prediction as a threshold and calculate the corresponding precision and recall and visualize this as a scatter plot. The histograms on the margins denote the number of such points within a bin. We observe that compared to Prompt-Naive, proposed method significantly improves the operational granularity by increasing the diversity and cardinality of the outputs, while outperforming it in terms of PRAUC. In addition, we observe that even though it is significantly more efficient and scalable, Ours-Sup-2call outperforms Sample-Prob while achieving comparable or higher density in the operational curves. Unlike the individual datasets, the performance of Ours-Unsup here is not comparable to that of the supervised ones. However, it still offers an easy way to increase the cardinality while outperforming Prompt-Naive. Figure 4 (bottom) shows PRAUC comparison for the joint dataset. Unlike on the individual datasets, the proposed supervised approaches and Sample-Prob vastly outperform other methods. With just $2$ samples per instance during inference, the Ours-Sup-2call achieves $8\%$ point improvement in mean PRAUC over the baseline Prompt-Naive and $2.6\%$ points over $20$ sample Sample-Prob approach. The learning approaches have relatively higher variance due to the small sizes of dataset-subsets that are merged to create the combined dataset. Table 3 shows the operational granularity measurements along $g^{pre}$, $g^{rec}$ and $g^{fpr}$, as well as the output cardinalities. Prompt-Naive and the two more complex prompting techniques Prompt-Seq and Prompt-Seq-CoT have an extremely small cardinality ($16$ or less). Sampling approaches improve it by nearly two orders of magnitude. However, the proposed methods achieve a cardinality nearly equaling the size of the dataset. This increase in cardinality also translates to better operating points, seen in the low granularity values along all $3$ axes of precision, recall and fpr, while maintaining or improving performance.
@@ 5 Conclusion
We studied the impact of prompting, confidence elicitation and uncertainty estimation techniques to black-box LLMs’ output distributions for decision making problems. We showed that verbalized probabilities have limited cardinality and provided possible explanations for why and when this issue surfaces. Our proposed approaches expand the range of operating points while maintaining or improving the overall prediction quality. Although empirically successful, our approaches do not directly optimize the objective in Equation 3, but rely on noise to increase granularity. This results in always operating in a granularity-performance tradeoff that may fail to retain performance on tasks with highly complex relationships between $\hat{y}$ and $y$. Future work includes moving away from this tradeoff by modeling $z$ as a prior to a generative model, and fitting a classifier to the representations of this model.